OpenStudy (anonymous):

what is the sum of an infinite geometric series with a first term of 6 and common ratio of 1/2

OpenStudy (anonymous):

Factor out the 6, then you have \[6(1 + 1/2 + 1/4 + ..)\]

OpenStudy (anonymous):

Do you know that one?

OpenStudy (anonymous):

ummm.. one sec

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

awesome! :D thanks man

OpenStudy (anonymous):

yw

OpenStudy (cwrw238):

formula is a / (1 - r) = 6 / ( 1 - 1/2) = 12